OpenStudy (anonymous):

furniture maker is making a chair and needs 12 feet of wood. He has chosen to make the chair out of maple, which costs $25 per foot, and pine, which costs $28 per foot. If the furniture maker wants to spend a total of $25.75 per foot, how many feet of each type of wood will he need?

OpenStudy (anonymous):

28x +25(12-x) = 12*25.75 28x + 300 - 25x = 309 3x = 9 x = 3 answer 3 ft of pine and 9 ft of maple